OpenStudy (anonymous):

Chris Dada can do a job in 9hr while Dana Bortes can do the same job in 5 hr. How long would it take them to do the job if they worked together?

OpenStudy (anonymous):

are you really typing

OpenStudy (anonymous):

This question is one of the most famous types of word problems; ones involving simultanious activities through time. Let's first examine Chris: Since it takes him 9 hours to do a job, therefore, in one hour, he can do 1/9th of the job, in two hours 2/9th of the job and so on. So in t hours he's do t/9 of the job. Looking at Dana, its the same thing, although she would be represented as t/5. So the equation would look something like this: t/9+t/5=1. The one would represent the entire job. So after cross multiplying, you should come out with an answer of 45/14 or 3.214 hours

OpenStudy (anonymous):

Would you like me to display the work?

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

please :)

OpenStudy (anonymous):

t/9+t/5=45 (total hours multiplyed together) (multiply out the 9 and 5 to get them out of the denominator) 45((t/5)+(t/9))=1(45) left with 9t+5t=45 9t+5t=14t 14t=45 divide both sides by 14 and you are left wit h45/14 or 3.2 hours Hope that helped lol :) -Top